25% of individuals in a certain city earns more than $45,000 per year.
The percentage of individuals earning more than $45,000 and having a college degree is 10. Suppose that a person is randomly chosen and turns out to be earning more than $45,000 . What is the probability that he or she is a college graduate.

Answer by stanbon(75887)

25% of individuals in a certain city earns more than $45,000 per year.
The percentage of individuals earning more than $45,000 and having a college degree is 10. Suppose that a person is randomly chosen and turns out to be earning more than $45,000 . What is the probability that he or she is a college graduate.
-------------
P(> 45000) = 0.25
P(> 45000 AND cd) = 0.10
P(cd |> 45000) = 0.10/0.25 = 2/5
------------------------
Cheers,
Stan H.
-----------